Archie Co. purchased a framing machine for $55,000 on January 1, 2021. The machine is expected to have a four-year life, with a residual value of $2,000 at the end of four years.

Using the sum-of-the years'-digits method, depreciation for 2022 and book value at December 31, 2022, would be: (Do not round intermediate calculations.)

Multiple Choice

  • $16,500 and $16,500 respectively.

  • $16,500 and $14,500 respectively.

  • $15,900 and $17,900 respectively.

  • $15,900 and $15,900 respectively.
Answer #1

Depreciation for 2021 = (55,000-2,000)*4/10 = 21,200

Depreciation for 2022 = (55,000-2,000)*3/10 = 15,900

Book value = 55,000 - (21,200+15,900) = 17,900

Option C is the answer